# Laura Barros        Aiko Schmeißer

### v3

@@ B. Candidacy
office, regardless of whether they are laid off or not. [^48] Interestingly, we observe relatively small differences between always-takers and never-takers on all measures of competence, suggesting that political selection may be weak, on average. For layoff-induced party members, we find similarly higher education, compared to always- and never-takers. Moreover, there are clear differences in log wages and AKM person effects (and their residuals) that also indicate higher competence among compliers. For instance, layoff-induced candidates have AKM person effects about 0.2 standard deviations higher than those of always- and never-takers, and for layoff-induced party members the difference is even larger.
@@ 6.2 Local-level evidence
In order to exploit plausibly exogenous changes in local labor demand, we follow Bartik (1991) and instrument $\Delta E_{mc}$ using variation in national industry employment growth rates interacted with baseline cross-sectional differences in municipalities’ industry composition. If industry employment shocks at the national level are exogenous, this strategy can be used to identify the effects of demand-driven variation in local employment (Borusyak et al., 2022). We construct the Bartik instruments as follows:
@@
where $s_{m,k,2004}$ is the initial employment share of industry $k$ in municipality $m$ and $e_{-m,k,c}$ is the national employment of industry $k$ leaving out municipality $m$ [^51] Industry- and municipality-level employment is constructed by counting all formal contracts in RAIS. We use 581 four-digit CNAE 1.0 industries.
(1) (2) (3) (4) Outcome: $\Delta$ Average years of education of … Councilor candidates Elected councilors [A] OLS $\Delta$ Employment rate (ppt) -0.035∗∗∗ -0.029∗∗∗ -$0∗∗$027∗∗∗ (0.003) (0.004) (0.006) (0.007) Observations 16,650 16,650 16,650 16,650 [B] 2SLS (Bartik IV) $\Delta$ Employment rate (ppt) -0.160∗∗∗ -0.150∗∗∗ -0.140∗∗∗ -0.122∗∗∗ (0.014) (0.016) (0.022) (0.025) KP F-stat 301.7 250.3 301.7 250.3 Observations 16,650 16,650 16,650 16,650 Fixed effects Year State $\times$ year Year State $\times$ year Note: The table presents OLS and 2SLS results from model (5) for the effect of formal employment rates on the average education of local councilors. The sample contains all municipal elections in 2004, 2008, 2012, and 2016. The dependent variable measures the change across subsequent elections in the average years of education of all individuals running (columns (1) and (2)) and being elected (columns (3) and (4)) for local council. The explanatory variable is the percentage point change in the municipal formal employment-to-population rate. For the 2SLS results, changes in employment rates are instrumented with Bartik instruments that are constructed by interacting baseline differences in municipalities’ industry composition with national industry employment growth rates (see equation (6)). All regressions are weighted by municipal population. Standard errors clustered at the municipality level are in parentheses. $p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$
@@
Appendix Tables D.3 and D.4 further show that negative local labor demands shocks can also improve political representation. Consistent with our individual-level analysis of compliers, we find that lower employment rates lead to an increase in the share of female and younger individuals who run for and are elected to councilor office. Together, these results highlight that negative economic shocks can enhance the competence of local politicians without compromising their socio-demographic representation.
@@ 7 Conclusion
The quality of democratic institutions largely depends on individuals’ willingness to engage in politics. While political careers plausibly compete with careers in the private sector, we lack causal evidence on how changes in labor market opportunities affect the trade-offs faced by individuals when deciding to enter politics. Combining rich information on the universe of party members and local candidates with matched employer-employee data covering all formal workers in Brazil, we study the effects of mass layoffs on individuals’ decisions to start a political career and the consequences of economic shocks for the selection of politicians.
@@ Appendix A Appendix to Sections 2 to 3
2004 2008 2012 2016 2020 All Total donations Mean 3,715.55 4,861.25 5,649.81 3,427.54 3,467.26 4,200.86 P10 202.69 211.58 212.38 230.70 214.83 213.61 P25 520.92 617.11 613.78 534.02 519.33 552.17 P50 1,403.30 1,833.70 2,123.80 1,468.55 1,327.48 1,586.85 P75 3,506.12 4,752.17 5,932.49 3,741.29 3,276.06 4,169.63 P90 7,994.55 10,655.85 13,199.48 7,435.77 7,450.96 9,300.73 Donations from own resources Share$>$0 0.61 0.68 0.55 0.56 0.46 0.56 Mean 1,632.70 1,718.96 1,704.98 1,225.80 624.07 1,313.14 P10 0.00 0.00 0.00 0.00 0.00 0.00 P25 0.00 0.00 0.00 0.00 0.00 0.00 P50 318.93 439.03 196.81 213.61 0.00 168.13 P75 1,811.53 1,819.59 1,792.28 1,281.64 681.86 1,275.73 P90 4,297.07 4,557.80 4,955.54 3,226.50 1,494.49 3,539.67 Observations 202,534 284,031 359,753 395,179 415,691 1,657,188 Note: The table shows summary statistics for campaign donations received by candidates for local council in the elections 2000 to 2020. We report total donation values and the value of donations that come from candidates’ own resources. All values are in BRL (CPI 2018) and are winsorized at the 1% level. The data is obtained from the Tribunal Superior Eleitoral.
@@ Appendix B Appendix to Section 4
  • Note: The table reports the average pre-layoff political outcomes of treated workers who are displaced from a mass layoff firm and matched control workers who are not displaced in the same calendar year, and the standardized difference between the two groups. The sample only includes workers with a unique name within the country. The ideological classification into ‘left’, ‘center’ and ‘right’ parties follows Colonnelli et al. (2022) and is shown in Appendix Table B.5. The ‘incumbent’ party is the party of the elected mayor in the individual’s municipality in 2008, and ‘not incumbent’ refers to all other parties. The ‘union-affiliated’ parties are PT, PDT, PSB, PCB, PSD, and MDB, and ‘not union-affiliated’ refer to all other left or center parties.
@@
A. Labor earnings B. Any employment Refer to caption Refer to caption C. Private-sector employment D. Public-sector employment Refer to caption Refer to caption Note: The figure reports event-study coefficients $\delta_{k}$, estimated in model (3), for the effect of job loss on yearly formal labor earnings (Panel A), the likelihood of being employed in any formal job at the end of the year (Panel B), and the likelihoods of being employed in the private sector (Panel C) and public sector (Panel D), separately for individuals with no degree, elementary school, or middle school ($<$ 12 years), high school (12 years), and university education ($>$ 12 years). The vertical bars depict 95% confidence intervals based on standard errors clustered at the individual level. Below each graph, the DiD coefficient from a static version of model (3), its standard error, and the mean of the control group across all post-treatment periods are reported.
Population Maximum wage Maximum wage in % of state in BRL, 2004 deputy salary 1-10,000 20% 1927.1 10,001-50,000 30% 2890.6 50,001-100,000 40% 3854.2 100,001-300,000 50% 4817.7 300,001-500,000 60% 5781.2 above 500,000 75% 7226.6 Note: The table shows the maximum monthly wage that councilors can earn by municipality population size. Wage caps are shown in % of state deputy salaries, as defined by the Constitutional Amendment No 25. in 2000, and in absolute values, as estimated by Ferraz and Finan (2011) based on the 2004 federal deputy salary.
@@
Left Center Right PT MDB PP PDT PSDB DEM PSB PTB PL CIDADANIA AVANTE PSC PCDOB PSD REPUBLICANOS PV PSL PMN PTC PSOL DC PCB PODE PSTU PRTB PCO PRP SD PHS PROS PATRIOTA PPL NOVO PMB REDE UP Note: The table shows the ideological classification of political parties, based on Colonnelli et al. (2022).
@@
A. New party membership B. Candidacy Refer to caption Refer to caption Note: The figure reports event-study coefficients $\delta_{c}$, estimated in model (1), for the effect of job loss on the likelihood of newly registering as a party member (Panel A) and running for local councilor (Panel B) in left, center, and right parties. The ideological classification of parties follows Colonnelli et al. (2022) and is shown Table B.5. The vertical bars depict 95% confidence intervals based on standard errors clustered at the individual level. Below each graph, the DiD coefficient from model (2), its standard error, and the mean of the control group in the post-treatment period (average across 2012, 2016, and 2020) are reported.
@@
Note: The figure reports DiD coefficients, estimated in model (2), for the effect of job loss on the likelihood of newly registering as a party member (Panel A) and running for local councilor (Panel B) in each of Brazil’s major parties. The ideological classification of parties follows Colonnelli et al. (2022). The horizontal bars depict 90% confidence intervals based on standard errors clustered at the individual level. All coefficients and standard errors are scaled by the mean of the control group in the post-treatment period (average across 2012, 2016, and 2020).
@@ Appendix C Appendix to Section 5
A. Layoff day in month B. Density around cutoff Refer to caption Refer to caption Note: Panel A reports a histogram for the distribution of layoff dates by calendar day in each month. Panel B reports the density of time between the current and the last layoff date around the 16-month cutoff date for eligibility for unemployment benefits, based on the initial sample that includes all layoff dates. The line represents a local linear polynomial smoothing using a 60-day bandwidth. Results for the bias-robust density test proposed by Cattaneo et al. (2018, 2020) are also shown.
@@
Refer to caption Note: The figure reports the density of time between the current and the last layoff date around the 16-month cutoff date for eligibility for unemployment benefits, based on our estimation sample that excludes all workers who were displaced on the first or the last day of each month in their last layoff. The line represents a local linear polynomial smoothing using a 60-day bandwidth. Results for the bias-robust density test proposed by Cattaneo et al. (2018, 2020) are also shown.
@@
(1) (2) (3) (4) (5) (6) (7) (8) (9) (10) Bandwidth (days): 30 opt 30 60 90 opt 90 120 150 opt Polynomial order: 0 0 1 1 1 1 2 2 2 2 [A] Full sample $\beta$ 0.092 0.097∗∗ 0.200∗∗ 0.150∗∗ 0.122∗∗ 0.162∗∗ 0.174 0.153∗∗ 0.150∗∗ 0.173 (0.051) (0.049) (0.102) (0.072) (0.059) (0.075) (0.089) (0.077) (0.069) (0.090) Control mean 3.500 3.493 3.463 3.467 3.479 3.454 3.465 3.461 3.458 3.450 Relative effect 2.6% 2.8% 5.8% 4.3% 3.5% 4.7% 5.0% 4.4% 4.3% 5.0% Observations 529,098 565,362 529,098 1,047,683 1,560,276 976,135 1,560,276 2,083,546 2,590,023 1,507,096 [B] Low educated ($<$ 12 years of education) $\beta$ 0.033 0.017 0.128 0.066 -0.031 0.091 0.100 0.008 0.016 0.112 (0.077) (0.074) (0.153) (0.109) (0.089) (0.112) (0.133) (0.115) (0.104) (0.136) Control mean 3.479 3.473 3.417 3.454 3.487 3.449 3.471 3.483 3.469 3.432 Relative effect 0.9% 0.5% 3.8% 1.9% -0.9% 2.6% 2.9% 0.2% 0.5% 3.3% Observations 229,191 244,890 229,191 454,665 682,272 430,963 682,272 922,954 1,157,173 650,326 [C] High educated ($\geq$ 12 years of education) $\beta$ 0.137∗∗ 0.137∗∗ 0.255 0.216∗∗ 0.244∗∗∗ 0.231∗∗ 0.230 0.262∗∗ 0.259∗∗∗ 0.214 (0.068) (0.068) (0.136) (0.097) (0.080) (0.094) (0.119) (0.104) (0.093) (0.121) Control mean 3.517 3.517 3.498 3.476 3.467 3.459 3.462 3.446 3.443 3.460 Relative effect 3.9% 3.9% 7.3% 6.2% 7.0% 6.7% 6.6% 7.6% 7.5% 6.2% Observations 299,907 299,907 299,907 593,018 878,004 641,120 878,004 1,160,592 1,432,850 848,505 Note: The table presents robustness checks for the effect of unemployment benefit eligibility on the likelihood of newly registering as a party member when using different bandwidths and polynomial orders of the running variable. ‘opt’ refers to the optimal bandwidth proposed by Calonico et al. (2014). Standard errors clustered at the individual level are in parentheses. The table also reports the control mean of the outcome at the cutoff and the effect sizes scaled by the control mean. All coefficients, standard errors, and control means have been scaled by 100, such that effects are interpreted in terms of percentage points. $p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$
@@
(1) (2) (3) (4) (5) (6) (7) (8) (9) (10) Bandwidth (days): 30 opt 30 60 90 opt 90 120 150 opt Polynomial order: 0 0 1 1 1 1 2 2 2 2 [A] Full sample $\beta$ 0.025 0.020 0.055∗∗ 0.028 0.030 0.030 0.031 0.024 0.025 0.030 (0.013) (0.013) (0.027) (0.019) (0.016) (0.020) (0.023) (0.020) (0.018) (0.024) Control mean 0.447 0.453 0.431 0.449 0.444 0.447 0.450 0.446 0.451 0.447 Relative effect 5.6% 4.3% 12.9% 6.2% 6.7% 6.7% 7.0% 5.3% 5.6% 6.6% Observations 1,035,237 1,105,868 1,035,237 2,048,674 3,058,455 1,908,388 3,058,455 4,101,059 5,107,090 2,953,238 [B] Low educated ($<$ 12 years of education) $\beta$ 0.001 -0.007 0.035 -0.006 -0.004 -0.003 -0.002 -0.008 -0.001 -0.004 (0.018) (0.018) (0.037) (0.026) (0.021) (0.027) (0.032) (0.028) (0.025) (0.033) Control mean 0.419 0.424 0.396 0.428 0.417 0.427 0.435 0.427 0.423 0.426 Relative effect 0.3% -1.6% 8.9% -1.4% -0.9% -0.6% -0.5% -1.9% -0.2% -1.0% Observations 497,269 531,075 497,269 986,755 1,485,874 934,928 1,485,874 2,021,245 2,540,083 1,414,964 [C] High educated ($\geq$ 12 years of education) $\beta$ 0.047∗∗ 0.047∗∗ 0.075∗∗ 0.061∗∗ 0.063∗∗∗ 0.072∗∗∗ 0.063 0.054 0.053∗∗ 0.063 (0.019) (0.019) (0.038) (0.027) (0.023) (0.026) (0.034) (0.029) (0.026) (0.034) Control mean 0.473 0.473 0.463 0.468 0.465 0.461 0.463 0.464 0.473 0.462 Relative effect 10.0% 10.0% 16.2% 13.0% 13.6% 15.6% 13.5% 11.6% 11.2% 13.6% Observations 537,968 537,968 537,968 1,061,919 1,572,581 1,148,350 1,572,581 2,079,814 2,567,007 1,519,603 Note: The table presents robustness checks for the effect of unemployment benefit eligibility on the likelihood of running for local councilor when using different bandwidths and polynomial orders of the running variable. ‘opt’ refers to the optimal bandwidth proposed by Calonico et al. (2014). Standard errors clustered at the individual level are in parentheses. The table also reports the control mean of the outcome at the cutoff and the effect sizes scaled by the control mean. All coefficients, standard errors, and control means have been scaled by 100, such that effects are interpreted in terms of percentage points. $p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$
@@
A. Full sample New party membership Candidacy Refer to caption Refer to caption B. Low educated ($<$ 12 years of education) New party membership Candidacy Refer to caption Refer to caption C. High educated ($\geq$ 12 years of education) New party membership Candidacy Refer to caption Refer to caption Note: The figures compare RD estimates for the effect of UI eligibility on the likelihood of newly registering as a party member and running for local councilor at the true 16-month cutoff (vertical red line) with the distribution of estimates at all possible placebo cutoffs within 180 days away from the true cutoff. The dashed (dotted) lines indicate the 2.5 (5) and 97.5 (95) percentiles of the distribution of placebo cutoffs. Estimates are based on the local linear regression model (4) with a bandwidth of 60 days. Panel A shows results among all workers in the sample, and Panels B and C distinguish workers with less than vs. at least 12 years of education.

### v4

@@ B. Candidacy
Figure 9 presents the results. Overall, we find that individuals who enter politics because of a layoff (compliers) tend to be positively selected in terms of competence relative to always-takers and never-takers. For example, layoff-induced councilor candidates have an average of 12.0 years of education, compared to about 10.9 years among individuals who always or never run for local office, regardless of whether they are laid off or not. [^51] Interestingly, we observe relatively small differences between always-takers and never-takers on all measures of competence, suggesting that political selection may be weak, on average. For layoff-induced party members, we find similarly higher education, compared to always- and never-takers. Moreover, there are clear differences in log wages and AKM person effects (and their residuals) that also indicate higher competence among compliers. For instance, layoff-induced candidates have AKM person effects about 0.2 standard deviations higher than those of always- and never-takers, and for layoff-induced party members the difference is even larger.
@@ 6.2 Local-level evidence
In order to exploit plausibly exogenous changes in local labor demand, we follow bartik1991 and instrument $\Delta E_{mc}$ using variation in national industry employment growth rates interacted with baseline cross-sectional differences in municipalities’ industry composition. If industry employment shocks at the national level are exogenous, this strategy can be used to identify the effects of demand-driven variation in local employment (borusyak2022). We construct the Bartik instruments as follows:
@@
where $s_{m,k,2004}$ is the initial employment share of industry $k$ in municipality $m$ and $e_{-m,k,c}$ is the national employment of industry $k$ leaving out municipality mm. [^54] Industry- and municipality-level employment is constructed by counting all formal contracts in RAIS. We use 581 four-digit CNAE 1.0 industries.
(1) (2) (3) (4) Outcome: $\Delta$ Average years of education of … Councilor candidates Elected councilors [A] OLS $\Delta$ Employment rate (ppt) -0.035∗∗∗ -0.029∗∗∗ -0.030∗∗∗ -0.027∗∗∗ (0.003) (0.004) (0.006) (0.007) Observations 16,650 16,650 16,650 16,650 [B] 2SLS (Bartik IV) $\Delta$ Employment rate (ppt) -0.160∗∗∗ -0.150∗∗∗ -0.140∗∗∗ -0.122∗∗∗ (0.014) (0.016) (0.022) (0.025) KP F-stat 301.7 250.3 301.7 250.3 Observations 16,650 16,650 16,650 16,650 Fixed effects Year State $\times$ year Year State $\times$ year Note: The table presents OLS and 2SLS results from model (5) for the effect of formal employment rates on the average education of local councilors. The sample contains all municipal elections in 2004, 2008, 2012, and 2016. The dependent variable measures the change across subsequent elections in the average years of education of all individuals running (columns (1) and (2)) and being elected (columns (3) and (4)) for local council. The explanatory variable is the percentage point change in the municipal formal employment-to-population rate. For the 2SLS results, changes in employment rates are instrumented with Bartik instruments that are constructed by interacting baseline differences in municipalities’ industry composition with national industry employment growth rates (see equation (6)). All regressions are weighted by municipal population. Standard errors clustered at the municipality level are in parentheses. $p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$
@@
Appendix Tables D.3 and D.4 further show that negative local labor demand shocks can also improve political representation. Consistent with our individual-level analysis of compliers, we find that lower employment rates lead to an increase in the share of female and younger individuals who run for and are elected to councilor office. Together, these results highlight that negative economic shocks can enhance the competence of local politicians without compromising their socio-demographic representation.
@@ 7 Conclusion
The quality of democratic institutions largely depends on individuals’ willingness to engage in politics. While political careers plausibly compete with careers in the private sector, we lack causal evidence on how changes in private labor market opportunities affect the trade-offs faced by individuals when deciding to enter politics. Combining rich information on the universe of party members and local candidates with matched employer-employee data covering all formal workers in Brazil, we study the effects of mass layoffs on individuals’ decisions to start a political career and the consequences of economic shocks for the selection of politicians.
@@ Appendix A Appendix to Sections 2 to 3
2004 2008 2012 2016 2020 All Total donations Mean 3,715.55 4,861.25 5,649.81 3,427.54 3,467.26 4,200.86 P10 202.69 211.58 212.38 230.70 214.83 213.61 P25 520.92 617.11 613.78 534.02 519.33 552.17 P50 1,403.30 1,833.70 2,123.80 1,468.55 1,327.48 1,586.85 P75 3,506.12 4,752.17 5,932.49 3,741.29 3,276.06 4,169.63 P90 7,994.55 10,655.85 13,199.48 7,435.77 7,450.96 9,300.73 Donations from own resources Share>>0 0.61 0.68 0.55 0.56 0.46 0.56 Mean 1,632.70 1,718.96 1,704.98 1,225.80 624.07 1,313.14 P10 0.00 0.00 0.00 0.00 0.00 0.00 P25 0.00 0.00 0.00 0.00 0.00 0.00 P50 318.93 439.03 196.81 213.61 0.00 168.13 P75 1,811.53 1,819.59 1,792.28 1,281.64 681.86 1,275.73 P90 4,297.07 4,557.80 4,955.54 3,226.50 1,494.49 3,539.67 Observations 202,534 284,031 359,753 395,179 415,691 1,657,188 Note: The table shows summary statistics for campaign donations received by candidates for local council in the elections 2000 to 2020. We report total donation values and the value of donations that come from candidates’ own resources. All values are in BRL (CPI 2018) and are winsorized at the 1% level. The data is obtained from the Tribunal Superior Eleitoral.
@@ Appendix B Appendix to Section 4
  • Note: The table reports the average pre-layoff political outcomes of treated workers who are displaced from a mass layoff firm and matched control workers who are not displaced in the same calendar year, and the standardized difference between the two groups. The sample only includes workers with a unique name within the country. The ideological classification into ‘left’, ‘center’ and ‘right’ parties follows Colonnelli2022 and is shown in Appendix Table B.5. The ‘incumbent’ party is the party of the elected mayor in the individual’s municipality in 2008, and ‘not incumbent’ refers to all other parties. The ‘union-affiliated’ parties are PT, PDT, PSB, PCB, PSD, and MDB, and ‘not union-affiliated’ refer to all other left or center parties.
@@
A. Labor earnings B. Any employment Refer to caption Refer to caption C. Private-sector employment D. Public-sector employment Refer to caption Refer to caption Note: The figure reports event-study coefficients $\delta_{k}$, estimated in model (3), for the effect of job loss on yearly formal labor earnings (Panel A), the likelihood of being employed in any formal job at the end of the year (Panel B), and the likelihoods of being employed in the private sector (Panel C) and public sector (Panel D), separately for individuals with no degree, elementary school, or middle school (<< 12 years), high school (12 years), and university education (>> 12 years). The vertical bars depict 95% confidence intervals based on standard errors clustered at the individual level. Below each graph, the DiD coefficient from a static version of model (3), its standard error, and the mean of the control group across all post-treatment periods are reported.
Population Maximum wage Maximum wage in % of state in BRL, 2004 deputy salary 1-10,000 20% 1927.1 10,001-50,000 30% 2890.6 50,001-100,000 40% 3854.2 100,001-300,000 50% 4817.7 300,001-500,000 60% 5781.2 above 500,000 75% 7226.6 Note: The table shows the maximum monthly wage that councilors can earn by municipality population size. Wage caps are shown in % of state deputy salaries, as defined by the Constitutional Amendment No 25. in 2000, and in absolute values, as estimated by Ferraz2011 based on the 2004 federal deputy salary.
@@
Left Center Right PT MDB PP PDT PSDB DEM PSB PTB PL CIDADANIA AVANTE PSC PCDOB PSD REPUBLICANOS PV PSL PMN PTC PSOL DC PCB PODE PSTU PRTB PCO PRP SD PHS PROS PATRIOTA PPL NOVO PMB REDE UP Note: The table shows the ideological classification of political parties, based on Colonnelli2022.
@@
A. New party membership B. Candidacy Refer to caption Refer to caption Note: The figure reports event-study coefficients $\delta_{c}$, estimated in model (1), for the effect of job loss on the likelihood of newly registering as a party member (Panel A) and running for local councilor (Panel B) in left, center, and right parties. The ideological classification of parties follows Colonnelli2022 and is shown Table B.5. The vertical bars depict 95% confidence intervals based on standard errors clustered at the individual level. Below each graph, the DiD coefficient from model (2), its standard error, and the mean of the control group in the post-treatment period (average across 2012, 2016, and 2020) are reported.
@@
Note: The figure reports DiD coefficients, estimated in model (2), for the effect of job loss on the likelihood of newly registering as a party member (Panel A) and running for local councilor (Panel B) in each of Brazil’s major parties. The ideological classification of parties follows Colonnelli2022. The horizontal bars depict 90% confidence intervals based on standard errors clustered at the individual level. All coefficients and standard errors are scaled by the mean of the control group in the post-treatment period (average across 2012, 2016, and 2020).
@@ Appendix C Appendix to Section 5
A. Layoff day in month B. Density around cutoff Refer to caption Refer to caption Note: Panel A reports a histogram for the distribution of layoff dates by calendar day in each month. Panel B reports the density of time between the current and the last layoff date around the 16-month cutoff date for eligibility for unemployment benefits, based on the initial sample that includes all layoff dates. The line represents a local linear polynomial smoothing using a 60-day bandwidth. Results for the bias-robust density test proposed by cattaneo2018; cattaneo2020 are also shown.
@@
Refer to caption Note: The figure reports the density of time between the current and the last layoff date around the 16-month cutoff date for eligibility for unemployment benefits, based on our estimation sample that excludes all workers who were displaced on the first or the last day of each month in their last layoff. The line represents a local linear polynomial smoothing using a 60-day bandwidth. Results for the bias-robust density test proposed by cattaneo2018; cattaneo2020 are also shown.
@@
(1) (2) (3) (4) (5) (6) (7) (8) (9) (10) Bandwidth (days): 30 opt 30 60 90 opt 90 120 150 opt Polynomial order: 0 0 1 1 1 1 2 2 2 2 [A] Full sample $\beta$ 0.092 0.097∗∗ 0.200∗∗ 0.150∗∗ 0.122∗∗ 0.162∗∗ 0.174 0.153∗∗ 0.150∗∗ 0.173 (0.051) (0.049) (0.102) (0.072) (0.059) (0.075) (0.089) (0.077) (0.069) (0.090) Control mean 3.500 3.493 3.463 3.467 3.479 3.454 3.465 3.461 3.458 3.450 Relative effect 2.6% 2.8% 5.8% 4.3% 3.5% 4.7% 5.0% 4.4% 4.3% 5.0% Observations 529,098 565,362 529,098 1,047,683 1,560,276 976,135 1,560,276 2,083,546 2,590,023 1,507,096 [B] Low educated (<< 12 years of education) $\beta$ 0.033 0.017 0.128 0.066 -0.031 0.091 0.100 0.008 0.016 0.112 (0.077) (0.074) (0.153) (0.109) (0.089) (0.112) (0.133) (0.115) (0.104) (0.136) Control mean 3.479 3.473 3.417 3.454 3.487 3.449 3.471 3.483 3.469 3.432 Relative effect 0.9% 0.5% 3.8% 1.9% -0.9% 2.6% 2.9% 0.2% 0.5% 3.3% Observations 229,191 244,890 229,191 454,665 682,272 430,963 682,272 922,954 1,157,173 650,326 [C] High educated ($\geq$ 12 years of education) $\beta$ 0.137∗∗ 0.137∗∗ 0.255 0.216∗∗ 0.244∗∗∗ 0.231∗∗ 0.230 0.262∗∗ 0.259∗∗∗ 0.214 (0.068) (0.068) (0.136) (0.097) (0.080) (0.094) (0.119) (0.104) (0.093) (0.121) Control mean 3.517 3.517 3.498 3.476 3.467 3.459 3.462 3.446 3.443 3.460 Relative effect 3.9% 3.9% 7.3% 6.2% 7.0% 6.7% 6.6% 7.6% 7.5% 6.2% Observations 299,907 299,907 299,907 593,018 878,004 641,120 878,004 1,160,592 1,432,850 848,505 Note: The table presents robustness checks for the effect of unemployment benefit eligibility on the likelihood of newly registering as a party member when using different bandwidths and polynomial orders of the running variable. ‘opt’ refers to the optimal bandwidth proposed by calonico2014. Standard errors clustered at the individual level are in parentheses. The table also reports the control mean of the outcome at the cutoff and the effect sizes scaled by the control mean. All coefficients, standard errors, and control means have been scaled by 100, such that effects are interpreted in terms of percentage points. $p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$
@@
(1) (2) (3) (4) (5) (6) (7) (8) (9) (10) Bandwidth (days): 30 opt 30 60 90 opt 90 120 150 opt Polynomial order: 0 0 1 1 1 1 2 2 2 2 [A] Full sample $\beta$ 0.025 0.020 0.055∗∗ 0.028 0.030 0.030 0.031 0.024 0.025 0.030 (0.013) (0.013) (0.027) (0.019) (0.016) (0.020) (0.023) (0.020) (0.018) (0.024) Control mean 0.447 0.453 0.431 0.449 0.444 0.447 0.450 0.446 0.451 0.447 Relative effect 5.6% 4.3% 12.9% 6.2% 6.7% 6.7% 7.0% 5.3% 5.6% 6.6% Observations 1,035,237 1,105,868 1,035,237 2,048,674 3,058,455 1,908,388 3,058,455 4,101,059 5,107,090 2,953,238 [B] Low educated (<< 12 years of education) $\beta$ 0.001 -0.007 0.035 -0.006 -0.004 -0.003 -0.002 -0.008 -0.001 -0.004 (0.018) (0.018) (0.037) (0.026) (0.021) (0.027) (0.032) (0.028) (0.025) (0.033) Control mean 0.419 0.424 0.396 0.428 0.417 0.427 0.435 0.427 0.423 0.426 Relative effect 0.3% -1.6% 8.9% -1.4% -0.9% -0.6% -0.5% -1.9% -0.2% -1.0% Observations 497,269 531,075 497,269 986,755 1,485,874 934,928 1,485,874 2,021,245 2,540,083 1,414,964 [C] High educated ($\geq$ 12 years of education) $\beta$ 0.047∗∗ 0.047∗∗ 0.075∗∗ 0.061∗∗ 0.063∗∗∗ 0.072∗∗∗ 0.063 0.054 0.053∗∗ 0.063 (0.019) (0.019) (0.038) (0.027) (0.023) (0.026) (0.034) (0.029) (0.026) (0.034) Control mean 0.473 0.473 0.463 0.468 0.465 0.461 0.463 0.464 0.473 0.462 Relative effect 10.0% 10.0% 16.2% 13.0% 13.6% 15.6% 13.5% 11.6% 11.2% 13.6% Observations 537,968 537,968 537,968 1,061,919 1,572,581 1,148,350 1,572,581 2,079,814 2,567,007 1,519,603 Note: The table presents robustness checks for the effect of unemployment benefit eligibility on the likelihood of running for local councilor when using different bandwidths and polynomial orders of the running variable. ‘opt’ refers to the optimal bandwidth proposed by calonico2014. Standard errors clustered at the individual level are in parentheses. The table also reports the control mean of the outcome at the cutoff and the effect sizes scaled by the control mean. All coefficients, standard errors, and control means have been scaled by 100, such that effects are interpreted in terms of percentage points. $p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$
@@
A. Full sample New party membership Candidacy Refer to caption Refer to caption B. Low educated (<< 12 years of education) New party membership Candidacy Refer to caption Refer to caption C. High educated ($\geq$ 12 years of education) New party membership Candidacy Refer to caption Refer to caption Note: The figures compare RD estimates for the effect of UI eligibility on the likelihood of newly registering as a party member and running for local councilor at the true 16-month cutoff (vertical red line) with the distribution of estimates at all possible placebo cutoffs within 180 days away from the true cutoff. The dashed (dotted) lines indicate the 2.5 (5) and 97.5 (95) percentiles of the distribution of placebo cutoffs. Estimates are based on the local linear regression model (4) with a bandwidth of 60 days. Panel A shows results among all workers in the sample, and Panels B and C distinguish workers with less than vs. at least 12 years of education.